OpenStudy (anonymous):

solve for a. 25a^2-4 = 12 Medal!!!

OpenStudy (haseeb96):

25a^2 -4=12 25a^2 =12 +4 25a^2 =16 a^2 = 16/25 now take square root on both sides a = 4 /5